Projectile Motion: Other applications


A boy throws a rock with speed y = 18.3 m/s at an angle of ∅ = 57.0
1
Expert's answer
2021-09-26T08:39:39-0400

Given:

"v_0=18.3\:\rm m/s"

"\theta=57.0^{\circ}"


The range

"R=\frac{v_0^2\sin2\theta}{g}=\frac{18.3^2\sin(114^{\circ})}{9.8}=31.2\:\rm m"

The maximim height

"h_{\max}=\frac{v_0^2\sin^2\theta}{2g}=\frac{18.3^2\sin^2(57.0^{\circ})}{2*9.8}=12.0\:\rm m"

The time of flight

"t=\frac{2v_0\sin\theta}{g}=\frac{2*18.3\sin57^{\circ}}{9.8}=3.13\:\rm s"